OpenStudy (anonymous):

Find the x intercepts, y=x^2-4x-5 are they (-5,0),(1,0)?

OpenStudy (bahrom7893):

x intercepts means y=0

OpenStudy (bahrom7893):

x^2-4x-5=0 (x-5)(x+1)=0

OpenStudy (bahrom7893):

So either x-5=0 x+1=0 Solve each of these to get your answer

OpenStudy (bahrom7893):

it's actually the opposite signs of those

OpenStudy (anonymous):

Ok so it would be x=5, and x=-1 i

OpenStudy (bahrom7893):

yup

OpenStudy (anonymous):

Did I type hem in the right way the first time? Just opposite numbers

OpenStudy (anonymous):

Awesome Thank you so much!! I really appreciate it!!

OpenStudy (bahrom7893):

well -5 is not 5, so that was wrong sorry, but the idea is to set y=0 when u are looking for x intercept and x=0 when looking for y-intercept

OpenStudy (anonymous):

ok so I should right my answer like, (5,0),(-1,0)

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

woo hooo thank you thank you!!!!